# Neural Internal Model Control: Learning a Robust Control Policy via Predictive Error Feedback

### v2

@@ IV-B Experiments on Quadrotors
As shown in Tab. II, our method consistently outperforms others across nearly all tasks, particularly under disturbances and out-of-distribution system dynamics. Compared to state-of-the-art learning-based methods, classical adaptive control methods perform worse due to their reliance on handcrafted parameters, which fail to generalize across diverse conditions. Among the baselines, DATT shows strong resilience to external disturbances, supported by its $\mathcal{L}_{1}$ estimator. However, its performance declines noticeably with unexpected dynamics in evaluation. Comparing DATT(GT) and DATT($\mathcal{L}_{1}$) reveals that the learned RL policy generalizes well when provided with ground-truth inputs, suggesting that the performance drop is primarily due to the handcrafted $\mathcal{L}_{1}$ estimator, which struggles to generalize across different systems. On the other hand. the teacher-student training approach in RMA results in the student policy underperforming compared to the teacher policy and shows instability under unseen systems, highlighting its unreliability for learning-based prediction.
In contrast, our method employs the proposed explicit model to establish a baseline prediction of the expected body transformation, effectively detecting discrepancies caused by internal and/or external sources. This informative feedback enhances the control policy’s responsiveness. Consequently, our method outperforms all baselines, including privileged variants like DATT(GT) and RMA(T), likely due to the informative predictive error feedback.

### v3

@@ Abstract
Index Terms:
Robust Control, Reinforcement Learning, Sensorimotor Learning
@@ IV-B Experiments on Quadrotors
As shown in Tab. II, our method consistently outperforms others across nearly all tasks, particularly under disturbances and out-of-distribution (OOD) system dynamics.
Among estimator-based baselines—both explicit (e.g., $\mathcal{L}_{1}$ estimator in $\mathcal{L}_{1}$-PID, $\mathcal{L}_{1}$-MPPI, DATT) and implicit (e.g., RMA’s neural encoder)—DATT is the most resilient, benefiting from its explicit $\mathcal{L}_{1}$ estimator coupled with an RL-based controller. Nevertheless, all estimator-based methods degrade sharply under our extreme out-of-distribution scenarios—where internal dynamics shift or sudden external forces occur—because their estimators become unreliable and control robustness is compromised. Estimator-free approaches such as CAL and PPO also fail to manage these varied disturbances. In contrast, our method computes the predictive error between expected and observed body transformations, directly signaling state drift and disturbance magnitude. A policy trained on this error maintains strong performance across both nominal and unexpected conditions.
We also acknowledge that limited hyperparameter sweeps in our experiments may contribute to the baselines’ degraded performance. To ensure a fair comparison, we include privileged variants (DATT (GT) and RMA (T)) that receive ground-truth states rather than estimates, representing an upper tuning bound. Even against these privileged methods, our predictive error-driven controller achieves superior and more stable control, supporting our hypothesis that explicit predictive error feedback offers a simple, generalizable path to robust performance.